# How to Evaluate Distributed Coordination Systems? – A Survey and Analysis

### v2

@@ Abstract
Coordination services and protocols are critical components of distributed systems and are essential for providing consistency, fault tolerance, and scalability. However, due to the lack of a standard benchmarking tool for distributed coordination services, coordination service developers/researchers either use a NoSQL standard benchmark and omit evaluating consistency, distribution, and fault tolerance; or create their own ad-hoc microbenchmarks and skip comparability with other services. In this study, we analyze and compare known and widely used distributed coordination services, their evaluations, and the tools used to benchmark those systems. We identify important requirements of distributed coordination service benchmarking, like the metrics and parameters that need to be evaluated and their evaluation setups and tools.
@@
Distributed coordination, benchmarking, evaluation, wide-area networks.
@@ 1 Introduction
The difficulties in dealing with low-level Paxos and its variants in application development led to systems like ZooKeeper [41], Chubby [61], Tango [14], and WanKeeper [12] which abstracted away distributed coordination and provided frequently used coordination primitives as a service to the application developers. Many distributed applications and especially high-demand distributed data stores such as Google’s Spanner [24], Yahoo!’s PNUTS [22], Apache’s Mesos [37] and Kafka [49], and Twitter’s Manhattan [80] were built on top of such coordination services, each requiring different levels of synchronization, consistency, and availability guarantees.
Whether it is the development of a distributed coordination protocol or service or the development of an application employing these protocols/services, one of the significant challenges the developers face is the lack of standard benchmarking tools that could provide a comprehensive evaluation of the coordination framework: in terms of its performance, availability, scalability, and consistency guarantees. Currently, most developers use ad-hoc limited microbenchmarks to evaluate only a fraction of the functionality and use customized metrics and techniques. This results in a lack of comparability between coordination mechanisms and may lead to unfair claims of advantages over the competition.
@@ 2 Evaluation of Current Practices
We have analyzed the existing distributed coordination systems under three main categories: (i) consensus algorithms; (ii) coordination services; and (iii) distributed applications.
(i) Consensus Algorithms. These are the low-level algorithms and protocols designed to provide the necessary primitives for distributed coordination. Most of the systems in this category are variations of the original Paxos protocol, with improvements either in performance, availability, or scalability. The systems we have studied under this category include Mencius [62], FPaxos [39], Multi-Paxos [19], Hybrid-Paxos [27], Egalitarian Paxos (E-Paxos) [65], $M^{2}$ Paxos [71], Bizur [38], ZAB [45], WPaxos [11], SwiftPaxos [73], Omni-Paxos [67], and Hydra [21].
(ii) Coordination Services. These are high-level services that form a level of abstraction to hide the difficulties in dealing with low-level consensus algorithms and provide the frequently used coordination primitives as a service to the application developers. The systems studied under this category include ZooKeeper [41], Tango [14], Calvin [77], WanKeeper [12], ZooNet [55], Boki [42], FlexLog [33], and SplitFT [60].
@@ 2.1 Topologies and Experimental Setups
Flat topology is used in multi-leader or leaderless systems that allow concurrent updates like Mencius and E-Paxos. Although their way of assigning a replica to be the leader of a request differs, the topology represents the architecture of both systems. Similarly, the star topology is used in single leader protocols whether the system relies on strong leadership like ZooKeeper or changes the leader based on the outstanding requests at replicas as in Hybrid-Paxos. Multi-Star topology is used in systems with multiple quorums, each in a star topology, while leaders of every quorum communicate with each other and form a flat topology together (no hierarchy). Systems with partitioned data and access locality may use Multi-Star topology for better read performance or improved fault tolerance as in ZooNet and $M^{2}$ Paxos. Systems with dynamic quorums according to their data partition and replication requirements as in Spanner may also use this topology. Hierarchical topology can be defined as a multi-star topology with a hierarchy between the leaders. It is used in WanKeeper mainly to control the distribution of ownership of data objects in wide-area networks, as the master is the only replica to change the ownership. FPaxos uses grid topology and grid quorum to improve the performance in a cluster, while WPaxos benefits from the same to decrease the wide-area networking (WAN) communication overhead for the write operations. Central-log topology is used by systems like Tango, Boki, and Calvin which keep the execution order of transactions in a durable, consistent, and fault-tolerant shared log.
[…]
@@
In Table 2.1, the experimental setups of the studied systems are given. The number of regions field shows the level of geo-distribution. If systems are deployed in real-world computing environments such as public clouds (as in ZooNet and WanKeeper) or geographically distributed private networks (as in PNUTS), and the round trip time (RTT) between distributed replicas is greater than 20ms, it is included as a region. Similarly, for evaluations in controlled environments, if RTT between modeled regions is larger than 20ms, as in Mencius and Hybrid-Paxos, it is also accepted as a separate region. However, if all the servers are located in the same placement group (as in $M^{2}$ Paxos) or clusters in close proximity (as in Spanner and Calvin), it is considered a single region deployment even if the system design parameters support wide-area network deployments. The number of servers field represents the level of replication. The source of the produced workload is defined as the client, and the total number of workload sources is given as the number of clients.
@@
Systems apply different ways of creating client tasks. The general approach is to utilize a separate machine to create the required workload. For single cluster evaluations using a separate machine, one or multiple threads are used as clients, as in Bizur and Multi-Paxos, respectively. In WAN deployments, clients are distributed to the regions uniformly, as in Mencius, E-Paxos, and WPaxos. On the other hand, ZAB, FPaxos, and DistibutedLog do not implement external clients. Similarly, $M^{2}$ Paxos uses separate threads on the replica servers to generate the workload. ZooKeeper (35 servers for 250 clients) and ZooNet (2 servers for 60 clients) combine these approaches and use separate machines with multiple client threads. We provide testbed information to clarify whether the evaluations were simulated under a controlled environment like Emulab [44] and DETER [16] or performed on real computing environments like CloudLab [28], Amazon EC2 [2], Alibaba ECS [1], and Google Compute Engines [7]. This information is important to understand the involvement of external factors in evaluations such as variance in routing overhead as it is fully deterministic in controlled environments while you have limited control in real computing environments.
@@
Custom testbeds can be considered in between as they are closer to controlled environments for single-region deployments and closer to cloud computing environments for geo-distributed deployments. Under the benchmark field, if a standard benchmark is not used for the evaluations, it is listed as ”Microbenchmark.” Different implementations in evaluations of these custom microbenchmarks are given under the related metrics. Table  2.1 lists all of the used experimental setups in the evaluations of studied systems together, although some of them change their experimental setup for different metrics. For instance, Mencius uses 3 servers in 3 regions for performance evaluations; however, it uses 3,5,7 servers in 3,5,7 regions for the scalability benchmarks. Similarly, Calvin studies conflicting command performance with 4 and 8 replicas while using 100 replicas to conduct scalability experiments. If the different experimental setups of the same system are important, they are explained in detail in the sections of related metrics. For any of the fields in the table, if the details are not given explicitly in the related published materials, they are marked as ”Not Specified” in the table.
@@ 2.2 Metrics
The majority of the protocols and systems we have studied are evaluated for several different metrics. These metrics describe various aspects of system behavior under the workload, such as performance, scalability, availability, and consistency. Despite the lack of unity in the benchmarking of the systems, many authors have chosen to use the same metrics for their evaluations. In the rest of this section, we will describe the benchmarking metrics used in the literature and how these metrics allow researchers to evaluate their systems. Table 2.2 summarizes our findings.
@@
Applications Spanner  [24]
@@
[…]
Access
[…]
Objects
Algorithms Mencius [62] 50 100 NS 16,128,1024 6B,4KB
@@
Applications Spanner  [24] 0,100 NS NS NS 4KB
@@
Systems follow different approaches for evaluating performance. Although the majority of the systems evaluate the throughput and latency together, some systems such as Calvin, ZooNet, Multi-Paxos, and Omni-Paxos measure only throughput, and systems such as ZooKeeper and Spanner evaluate the throughput and latency separately. Evaluating latency and throughput together is essential to understanding the system’s overall performance. While the throughput is mainly related to the system’s processing capacity, latency is more affected by I/O operations [30].
The performance of any distributed coordination system highly depends on workload characteristics such as read/write ratio, data access overlap, access locality, the number of data objects (size of data pool), and the size of data objects. In table 2.2, we list how the systems tuned their workloads for these parameters.
@@
Read/Write Ratio is a fundamental parameter to show the system’s performance under different use case scenarios. Since many systems handle write and read requests differently and the percentage of write and read requests is application-specific, it plays a crucial role in showing the applicability of the systems for various purposes. For instance, systems designed to be used for different aspects of distributed coordination like ZooKeeper or WanKeeper show the performance evaluation continuously while varying the write ratio from 0 to 100%. Similarly, Spanner measures the read and write performance separately since read requests can be handled by replicas, but writes need to be done by the leader. Besides being handled by the leader, writes require replication, which increases commit latency. To measure the true performance of replication, Multi-Paxos also uses 100% write operations. Due to the differences in processing read requests between systems used for comparison in the evaluations, Bizur uses only write operations. E-Paxos and Hybrid Paxos also use 100% writes since they focus on handling conflicting commands, and read requests do not produce conflicts. Some works like Calvin, CockroachDB, and OceanBase do not explicitly provide a write ratio for their evaluations but use standard benchmarks like TPC-C [20] and YCSB [23] with known write ratios.
@@
Data access overlap may have a major impact on the throughput and latency in multi-leader systems and protocols. It can be defined as the percentage of the key space shared by all clients. In this matter, data access overlap is 100% if the entire key space is shared and all clients are allowed to access any key. On the other hand, data access overlap is 0% if the key space is partitioned for each client and clients are accessing only the keys in their partition. For single leader algorithms providing strong consistency like Multi-Paxos, ZAB, Chubby, and ZooKeeper varying data access overlap has no impact on the overall performance since all write requests are treated in the same way. Data access overlap is an important factor for the performance of multi-leader algorithms/systems that allow concurrent update operations. Receiving different execution orders at different replicas is considered a command conflict. A command conflict is usually resolved by serializing the requests from different clients to have the same execution order or by checking the dependency between the requests to make sure the order of execution does not lead to different final states of the system. Mencius and E-Paxos can be considered under this category and they evaluate their system under the worst condition with 100% data access overlap. Hybrid-Paxos is also considered to be evaluated with 100% data access overlap since all clients are updating the same data object (one bank account). For these systems, data access overlap can also result in command collisions, as described in 2.2, which occurs when multiple servers update the same data object concurrently. Systems that use ownership to control the access to data objects are also affected by conflict ratio due to ownership migration or remote requests. Systems using static ownership, like ZooNet, forward all requests to the remote server, which is possibly located in a geographically far region, and they are affected by network delays for the requests of a not-owned data object. The use of buckets in Bizur is similar. Systems with dynamic ownership are either affected by exhibiting longer latency for remote requests or performing costly ownership migration, as in $M^{2}$ Paxos, WPaxos, WanKeeper, and PNUTS.
@@
Access locality is defined as the likelihood of clients accessing a specific part of the key space. If there is a 70% access locality, it represents the case that 70% of the requests are related to the same part of the key space, possibly the same region, and the other 30% of the requests are related to the rest of the key space. Access locality is more of a matter of the distribution of client requests rather than the sharing of the key space among clients which is defined as data access overlap. Clients may have different access localities while preserving 100% data access overlap as in the above example. This parameter has a big impact on performance in systems that use object ownership to parallelize request execution.
Access locality may not necessarily correspond to geographical partitioning of data, and in more general terms, it can be seen as the probability of accessing some preferred subset of keys, as illustrated in $M^{2}$ Paxos. WPaxos and COPS adjust the access locality by distributing the data objects uniformly in data groups and changing the client access rate for the data groups. While COPS changes the access locality for each client from 100% to 0%, WPaxos evaluates the access locality for 70% and 90%. $M^{2}$ Paxos measures the performance for the worst and best conditions as 0% and 100% access locality. ZooNet and PNUTS distribute the ownership of the data objects to adjust the level of local execution so their access locality rates correspond to the rate of local execution. Similarly, Calvin partitions its data to multiple datacenters and multiple machines in the same datacenter. Calvin’s evaluations exhibit a 90% access locality since it accesses the data objects stored at the same machine for 90% of the cases. PNUTS also examines the effect of non-uniform access on data objects by using Zipfian distribution [9] with varying Zipf factors which results in different access locality rates.
@@
The size of objects, as another parameter of performance evaluation, is usually used as a factor of saturation type. To make the servers CPU-bound, the size is kept as small as possible, and for the network-bound workloads, system evaluation may use larger objects. It is also related to the data models and the expected use of the systems. ZooKeeper, Tango, Spanner, Boki, FlexLog, and ZooNet use 1KB or 4 KB sized data objects as these are more suitable to their data model. Mencius and Multi-Paxos evaluate the effect of the data object size to evaluate the system for both CPU and network-bound cases.
@@
The ability of any system to perform well in a wide range of workload parameters and configurations is essential. The scalability measured by the system is of two distinct types: workload scalability and system scalability. Both types are often evaluated by measuring the changes in performance as some parameters are controlled. The parameters important for scalability evaluation are often those that describe the experimental testbed and can be seen in Table 2.1.
@@
To evaluate workload scalability, a certain amount of workload should be created until systems reach their saturation point. The amount of workload can easily be adjusted by either changing the number of clients or the number of requests per client. Hybrid-Paxos, E-Paxos, Tango, ScalarDB, and PNUTS increase the number of concurrent clients with a certain amount of workload produced by each client for the workload scalability evaluations. $M^{2}$ Paxos also increases the number of clients but it is more related to producing a new workload for the newly added server. Thus, it is not considered directly related to workload scalability evaluation. In some cases, clients may be throttled down, and changing the degree of throttling can be used to control the workload intensity, as it is done in WPaxos.
@@
Depending on the underlying algorithms, systems differ in how they scale with the increasing number of servers/replicas. Paxos derivations generally provide low performance for a higher number of replicas. Some systems, such as ZooKeeper, however, can provide increased read throughput as the number of replicas grows since these systems do not put a single leader in a read path and allow reading directly from replicas. Some works like Eris [57], NOPaxos [58], and HydraPaxos [21] sidestep consensus with in-network sequencing and groupcasts to achieve better performance with more replicas than consensus-based derivations. Increasing the number of servers/replicas does not always increase the aggregated throughput linearly due to the communication and replication overheads for most systems. To evaluate system scalability as the number of servers/replicas increases, evaluations often resort to CPU-bound workloads, as was done by $M^{2}$ Paxos, ZooKeeper, Tango, Calvin, Spanner, and E-Paxos. The differences in the system architecture and the underlying algorithms are also reflected in the methodologies for the system scalability experiments. $M^{2}$ Paxos used 100% access locality to eliminate the overhead due to ownership migration and tested for up to 49 servers while ZooKeeper uses up to 13 replicas for varying the write ratio from 0% to 100%. On the other hand, Tango evaluates for up to 100 servers since its scalability heavily relies on the underlying shared log, CORFU [13].

### v3

@@ Abstract
Coordination services and protocols are critical components of distributed systems and are essential for providing consistency, fault tolerance, and scalability. However, due to the lack of standard benchmarking and evaluation tools for distributed coordination services, coordination service developers/researchers either use a NoSQL standard benchmark and omit evaluating consistency, distribution, and fault tolerance; or create their own ad-hoc microbenchmarks and skip comparability with other services. In this study, we analyze and compare the evaluation mechanisms for known and widely used consensus algorithms, distributed coordination services, and distributed applications built on top of these services. We identify the most important requirements of distributed coordination service benchmarking, such as the metrics and parameters for the evaluation of the performance, scalability, availability, and consistency of these systems. Finally, we discuss why the existing benchmarks fail to address the complex requirements of distributed coordination system evaluation.
@@
Distributed coordination, benchmarking, scalability, availability, consistency, evaluation mechanisms.
@@ 1 Introduction
The difficulties in dealing with low-level Paxos and its variants in application development led to systems like ZooKeeper [43], Chubby [63], Tango [15], and WanKeeper [12], which abstracted away distributed coordination and provided frequently used coordination primitives as a service to the application developers. Many distributed applications and especially high-demand distributed data stores such as Google’s Spanner [25], Yahoo!’s PNUTS [23], Apache’s Mesos [39] and Kafka [51], and Twitter’s Manhattan [82] were built on top of such coordination services, each requiring different levels of synchronization, consistency, and availability guarantees.
Whether it is the development of a distributed coordination protocol or service or the development of an application employing these protocols/services, one of the significant challenges the developers face is the lack of standard benchmarking tools that could provide a comprehensive evaluation of the coordination framework – in terms of its performance, availability, scalability, and consistency guarantees. Currently, most developers use ad-hoc limited microbenchmarks to evaluate only a fraction of the functionality and use customized metrics and techniques. This results in a lack of comparability between coordination mechanisms and may lead to unfair claims of advantages over the competition.
@@ 2 Evaluation of Current Practices
We have analyzed the existing distributed coordination systems under three main categories: (i) consensus algorithms; (ii) coordination services; and (iii) distributed applications. The relevance, recency, and impact of the systems were considered during the selection process.
(i) Consensus Algorithms. These are the low-level algorithms and protocols designed to provide the necessary primitives for distributed coordination. Most of the systems in this category are variations of the original Paxos protocol, with improvements either in performance, availability, or scalability. The systems we have studied under this category include Mencius [64], FPaxos [41], Raft [70], Multi-Paxos [20], Hybrid-Paxos [29], Egalitarian Paxos (E-Paxos) [67], $M^{2}$ Paxos [73], Bizur [40], ZAB [47], WPaxos [11], SwiftPaxos [75], Omni-Paxos [69], and Hydra [22].
(ii) Coordination Services. These are high-level services that form a level of abstraction to hide the difficulties in dealing with low-level consensus algorithms and provide the frequently used coordination primitives as a service to the application developers. The systems studied under this category include ZooKeeper [43], Tango [15], Calvin [79], WanKeeper [12], ZooNet [57], Boki [44], FlexLog [35], SplitFT [62], Fabric [13], and Narwhal [26].
@@ 2.1 Topologies and Experimental Setups
Flat topology is used in multi-leader or leaderless systems that allow concurrent updates like Mencius and E-Paxos. Although their way of assigning a replica to be the leader of a request differs, the topology represents the architecture of both systems. Similarly, the star topology is used in single leader protocols, whether the system relies on strong leadership like ZooKeeper or changes the leader based on the outstanding requests at replicas, as in Hybrid-Paxos. Multi-Star topology is used in systems with multiple quorums, each in a star topology, while leaders of every quorum communicate with each other and form a flat topology together (no hierarchy). Systems with partitioned data and access locality may use Multi-Star topology for better read performance or improved fault tolerance as in ZooNet and $M^{2}$ Paxos. Systems with dynamic quorums, according to their data partition and replication requirements, as in Spanner, may also use this topology. Hierarchical topology can be defined as a multi-star topology with a hierarchy between the leaders. It is used in WanKeeper mainly to control the distribution of ownership of data objects in wide-area networks, as the master is the only replica to change the ownership. FPaxos uses grid topology and grid quorum to improve the performance in a cluster, while WPaxos benefits from the same to decrease the wide-area networking (WAN) communication overhead for the write operations. Central-log topology is used by systems like Tango, Boki, and Calvin, which keep the execution order of transactions in a durable, consistent, and fault-tolerant shared log.
System     # ofRegions     # ofServers     # ofClients Topology Testbed Benchmark
Algorithms Mencius [64] 3,5,7 3,5,7 3,5,7 Flat DETER Microbenchmark
FPaxos  [41] 1 5,8 Grid Custom Microbenchmark
Raft  [70] 1 5,9 Star Custom Microbenchmark
Multi-Paxos  [20] 1 5 1,4,10,20 Star Custom Microbenchmark
Hybrid Paxos  [29] 5,7,11,21 5,7,11,21 10,20,100,1000 Star Emulab Microbenchmark
E-Paxos [67] 1,3,5 3,5 50 Flat EC2 Microbenchmark
$M^{2}$ Paxos  [73] 1 5,11,49 64/server Multi-Star EC2 Microbenchmark
Bizur  [40] 1 3 1 Flat Custom Microbenchmark
ZAB  [47] 1 3-13 Star Custom Microbenchmark
WPaxos  [11] 5 15 5 Grid EC2 Microbenchmark
SwiftPaxos  [75] 13 5 1000,5000 Star EC2 YCSB  [24]
Omni-Paxos  [69] 1,3 3,5 1 Flat Google C.E. Microbenchmark
Hydra  [22] 1 8 StarHierarchical Custom MicrobenchmarkYCSB+T [28]
      CoordinationServices ZooKeeper  [43] 1 3-13 250 Star Custom Microbenchmark
Tango  [15] 1 18 18 Central-Log Custom YCSB  [24]
Calvin  [79] 1 4,8,0-100 Central-Log EC2 TPC-C  [8]
WanKeeper  [12] 3 3 1,2 Hierarchical EC2 YCSB  [24]
ZooNet  [57] 2 8 60 Multi-Star Google C.E. Microbenchmark
Boki  [44] 1 11, 64 64, 96, 128, 192 Central-Log EC2 MicrobenchmarkDeathStar [34]
FlexLog  [35] 1 1-6 1 Central-Log Custom Microbenchmark
SplitFT  [62] 1 1, 20 5, 43 Star CloudLab MicrobenchmarkYCSB [24]
Fabric  [13] 1, 2, 5 15-110 Flat IBM Cloud Microbenchmark
Narwhal  [26] 5 8—51 1—10 Flat EC2 Microbenchmark
Applications Spanner  [25] 1 1,3,5 100 Multi-Star Custom Microbenchmark
DistibutedLog  [37] 2 20 Star Custom Microbenchmark
PNUTS  [23] 3 3 300 Multi-Star Custom Microbenchmark
COPS  [61] 1 2,4,8,16,32 2,1024/server Multi-Star VICCI Microbenchmark
CockroachDB [5] 1 3,30 Flat Custom            TPC-C [8]Jepsen [49]
OceanBase  [85] 3 1557 360000/server Flat Alibaba ECS TPC-C [8]
ScalarDB  [84] 1 2, 203-1015 4, 8 Star EC2 Elle [50]YCSB [24]TPC-C [8]
@@
In Table I, the experimental setups of the studied systems are given. The number of regions field shows the level of geo-distribution. If systems are deployed in real-world computing environments such as public clouds (as in ZooNet and WanKeeper) or geographically distributed private networks (as in PNUTS), and the round-trip time (RTT) between distributed replicas is greater than 20ms, it is included as a region. Similarly, for evaluations in controlled environments, if RTT between modeled regions is larger than 20ms, as in Mencius and Hybrid-Paxos, it is also accepted as a separate region. However, if all the servers are located in the same placement group (as in $M^{2}$ Paxos) or clusters in close proximity (as in Spanner and Calvin), it is considered a single region deployment even if the system design parameters support wide-area network deployments. The number of server fields represents the level of replication. The source of the produced workload is defined as the client, and the total number of workload sources is given as the number of clients.
@@
Systems apply different ways of creating client tasks. The general approach is to utilize a separate machine to create the required workload. For single-cluster evaluations using a separate machine, one or multiple threads are used as clients, as in Bizur and Multi-Paxos, respectively. In WAN deployments, clients are distributed to the regions uniformly, as in Mencius, E-Paxos, and WPaxos. On the other hand, ZAB, FPaxos, and DistributedLog do not implement external clients. Similarly, $M^{2}$ Paxos uses separate threads on the replica servers to generate the workload. ZooKeeper (35 servers for 250 clients) and ZooNet (2 servers for 60 clients) combine these approaches and use separate machines with multiple client threads. We provide testbed information to clarify whether the evaluations were simulated under a controlled environment like Emulab [46] and DETER [17] or performed on real computing environments like CloudLab [30], Amazon EC2 [2], Alibaba ECS [1], and Google Compute Engines [7]. This information is important to understand the involvement of external factors in evaluations, such as variance in routing overhead, as it is fully deterministic in controlled environments, while you have limited control in real computing environments.
@@
Custom testbeds can be considered in between as they are closer to controlled environments for single-region deployments and closer to cloud computing environments for geo-distributed deployments. Under the benchmark field, if a standard benchmark is not used for the evaluations, it is listed as ”Microbenchmark.” Different implementations in evaluations of these custom microbenchmarks are given under the related metrics. Table I lists all of the experimental setups used in the evaluations of studied systems together, although some of them change their experimental setup for different metrics. For instance, Mencius uses 3 servers in 3 regions for performance evaluations; however, it uses 3,5,7 servers in 3,5,7 regions for the scalability benchmarks. Similarly, Calvin studies conflicting command performance with 4 and 8 replicas while using 100 replicas to conduct scalability experiments. If different experimental setups of the same system are important, they are explained in detail in the section of the corresponding metric. For any of the fields in the table, if the details are not given explicitly in the related published materials, they are marked as ”Not Specified” in the table.
@@ 2.2 Metrics
The majority of the protocols and systems we have studied are evaluated for several different metrics. These metrics describe various aspects of system behavior under the workload, such as performance, scalability, availability, and consistency. Despite the lack of unity in the benchmarking of the systems, many authors have chosen to use the same metrics for their evaluations. In the rest of this section, we will describe the benchmarking metrics used in the literature and how these metrics allow researchers to evaluate their systems. Table II summarizes our findings.
@@
Fabric [13]
Narwhal [26]
Applications
Spanner  [25]
@@
System      WriteRatio %       Data     AccessOverlap %           AccessLocality %    # ofObjects     Size ofObjects
Algorithms
Mencius [64] 50 100 NS 16,128,1024 6B,4KB
@@
Fabric  [13] 50-100 100 NS $2^{19}$ 0.5MB-4MB
Narwhal  [26] 100 NS NS NS 0.5MB
Applications
Spanner  [25] 0,100 NS NS NS 4KB
@@
Systems follow different approaches for evaluating performance. Although the majority of the systems evaluate the throughput and latency together, some systems, such as Calvin, ZooNet, Multi-Paxos, and Omni-Paxos, measure only throughput, and systems such as ZooKeeper and Spanner evaluate the throughput and latency separately. Evaluating latency and throughput together is essential to understanding the system’s overall performance. While the throughput is mainly related to the system’s processing capacity, latency is more affected by I/O operations [32].
The performance of any distributed coordination system highly depends on workload characteristics such as read/write ratio, data access overlap, access locality, the number of data objects (size of data pool), and the size of data objects. In table III, we list how the systems tuned their workloads for these parameters.
@@
Read/Write Ratio is a fundamental parameter to show the system’s performance under different use case scenarios. Since many systems handle write and read requests differently, and the percentage of write and read requests is application-specific, it plays a crucial role in showing the applicability of the systems for various purposes. For instance, systems designed to be used for different aspects of distributed coordination, like ZooKeeper or WanKeeper, show the performance evaluation continuously while varying the write ratio from 0 to 100%. Similarly, Spanner measures the read and write performance separately since read requests can be handled by replicas, but writes need to be done by the leader. Besides being handled by the leader, writes require replication, which increases commit latency. To measure the true performance of replication, Multi-Paxos also uses 100% write operations. Due to the differences in processing read requests between systems used for comparison in the evaluations, Bizur uses only write operations. E-Paxos and Hybrid Paxos also use 100% writes since they focus on handling conflicting commands, and read requests do not produce conflicts. Some works like Calvin, CockroachDB, and OceanBase do not explicitly provide a write ratio for their evaluations but use standard benchmarks like TPC-C [21] and YCSB [24] with known write ratios.
@@
Data access overlap may have a major impact on the throughput and latency in multi-leader systems and protocols. It can be defined as the percentage of the key space shared by all clients. In this matter, data access overlap is 100% if the entire key space is shared and all clients are allowed to access any key. On the other hand, data access overlap is 0% if the key space is partitioned for each client and clients are accessing only the keys in their partition. For single leader algorithms providing strong consistency like Multi-Paxos, ZAB, Chubby, and ZooKeeper, varying data access overlap has no impact on the overall performance since all write requests are treated in the same way. Data access overlap is an important factor for the performance of multi-leader algorithms/systems that allow concurrent update operations. Receiving different execution orders at different replicas is considered a command conflict. A command conflict is usually resolved by serializing the requests from different clients to have the same execution order or by checking the dependency between the requests to make sure the order of execution does not lead to different final states of the system. Mencius and E-Paxos can be considered under this category and they evaluate their system under the worst condition with 100% data access overlap. Hybrid-Paxos is also considered to be evaluated with 100% data access overlap since all clients are updating the same data object (one bank account). For these systems, data access overlap can also result in command collisions, as described in a, which occurs when multiple servers update the same data object concurrently. Systems that use ownership to control access to data objects are also affected by the conflict ratio due to ownership migration or remote requests. Systems using static ownership, like ZooNet, forward all requests to the remote server, which is possibly located in a geographically far region, and they are affected by network delays for requests of a non-owned data object. The use of buckets in Bizur is similar. Systems with dynamic ownership are either affected by exhibiting longer latency for remote requests or performing costly ownership migration, as in $M^{2}$ Paxos, WPaxos, WanKeeper, and PNUTS.
@@
Access locality is defined as the likelihood of clients accessing a specific part of the key space. If there is a 70% access locality, it represents the case that 70% of the requests are related to the same part of the key space, possibly the same region, and the other 30% of the requests are related to the rest of the key space. Access locality is more of a matter of the distribution of client requests rather than the sharing of the key space among clients, which is defined as data access overlap. Clients may have different access localities while preserving 100% data access overlap as in the above example. This parameter has a big impact on performance in systems that use object ownership to parallelize request execution.
Access locality may not necessarily correspond to geographical partitioning of data, and in more general terms, it can be seen as the probability of accessing some preferred subset of keys, as illustrated in $M^{2}$ Paxos. WPaxos and COPS adjust the access locality by distributing the data objects uniformly in data groups and changing the client access rate for the data groups. While COPS changes the access locality for each client from 100% to 0%, WPaxos evaluates the access locality for 70% and 90%. $M^{2}$ Paxos measures the performance for the worst and best conditions as 0% and 100% access locality. ZooNet and PNUTS distribute the ownership of the data objects to adjust the level of local execution so their access locality rates correspond to the rate of local execution. Similarly, Calvin partitions its data to multiple datacenters and multiple machines in the same datacenter. Calvin’s evaluations exhibit a 90% access locality since it accesses the data objects stored on the same machine for 90% of the cases. PNUTS also examines the effect of non-uniform access on data objects by using Zipfian distribution [9] with varying Zipf factors, which results in different access locality rates.
@@
The size of objects, as another parameter of performance evaluation, is usually used as a factor of the saturation type. To make the workload CPU-bound, the size is kept as small as possible, and for the network-bound workloads, system evaluation may use larger objects. It is also related to the data models and the expected use of the systems. ZooKeeper, Tango, Spanner, Boki, FlexLog, and ZooNet use 1KB or 4KB-sized data objects, as these are more suitable for their data model. Mencius and Multi-Paxos evaluate the effect of the data object size to evaluate the system for both CPU and network-bound cases.
@@
The ability of any system to perform well in a wide range of workload parameters and configurations is essential. The scalability measured by the system is of two distinct types: workload scalability and system scalability. Both types are often evaluated by measuring the changes in performance as some parameters are controlled. The parameters important for scalability evaluation are often those that describe the experimental testbed and can be seen in Table I.
@@
To evaluate workload scalability, a certain amount of workload should be created until systems reach their saturation point. The amount of workload can easily be adjusted by either changing the number of clients or the number of requests per client. Hybrid-Paxos, E-Paxos, Tango, ScalarDB, and PNUTS increase the number of concurrent clients with a certain amount of workload produced by each client for the workload scalability evaluations. $M^{2}$ Paxos also increases the number of clients, but it is more related to producing a new workload for the newly added server. Thus, it is not considered directly related to workload scalability evaluation. In some cases, clients may be throttled down, and changing the degree of throttling can be used to control the workload intensity, as it is done in WPaxos.
@@
Depending on the underlying algorithms, systems differ in how they scale with the increasing number of servers/replicas. Paxos derivations generally provide low performance for a higher number of replicas. Some systems, such as ZooKeeper, however, can provide increased read throughput as the number of replicas grows since these systems do not put a single leader in a read path and allow reading directly from replicas. Some works like Eris [59], NOPaxos [60], and HydraPaxos [22] sidestep consensus with in-network sequencing and groupcasts to achieve better performance with more replicas than consensus-based derivations. Increasing the number of servers/replicas does not always increase the aggregated throughput linearly due to the communication and replication overheads for most systems. To evaluate system scalability as the number of servers/replicas increases, evaluations often resort to CPU-bound workloads, as was done by $M^{2}$ Paxos, ZooKeeper, Tango, Calvin, Spanner, and E-Paxos. The differences in the system architecture and the underlying algorithms are also reflected in the methodologies for the system scalability experiments. $M^{2}$ Paxos used 100% access locality to eliminate the overhead due to ownership migration and tested for up to 49 servers, while ZooKeeper uses up to 13 replicas for varying the write ratio from 0% to 100%. On the other hand, Tango evaluates for up to 100 servers since its scalability heavily relies on the underlying shared log, CORFU [14].